### v1

# ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Performance Prediction

Abstract
In the realm of financial analytics, leveraging unstructured data, such as earnings conference calls (ECCs), to forecast stock performance is a critical challenge that has attracted both academics and investors. While previous studies have used deep learning-based models to obtain a general view of ECCs, they often fail to capture detailed, complex information. Our study introduces a novel framework: ECC Analyzer, combining Large Language Models (LLMs) and multi-modal techniques to extract richer, more predictive insights. The model begins by summarizing the transcript’s structure and analyzing the speakers’ mode and confidence level by detecting variations in tone and pitch for audio. This analysis helps investors form an overview perception of the ECCs. Moreover, this model uses the Retrieval-Augmented Generation (RAG) based methods to meticulously extract the focuses that have a significant impact on stock performance from an expert’s perspective, providing a more targeted analysis. The model goes a step further by enriching these extracted focuses with additional layers of analysis, such as sentiment and audio segment features. By integrating these insights, the ECC Analyzer performs multi-task predictions of stock performance, including volatility, value-at-risk (VaR), and return for different intervals. The results show that our model outperforms traditional analytic benchmarks, confirming the effectiveness of using advanced LLM techniques in financial analytics.
1 Introduction
The integration of structured data such as stock prices and financial ratios with unstructured data including financial filings and company news is increasingly essential in investment decision-making Fang and Zhang (2016); Roeder et al. (2022). This trend stems from the recognition that unstructured data can provide insights not fully captured by structured data alone. For instance, financial reports elucidate the broader implications of numerical data through discussions of managerial decisions and corporate strategies Wang and Hua (2014), while company news can shed light on public sentiment, emerging market trends, and market perceptions Kogan et al. (2009); Tetlock (2007). These nuanced insights are vital for a comprehensive analysis of complex market dynamics.
Advancements in Natural Language Processing (NLP) have significantly enhanced the ability to analyze unstructured data within the financial sector. Traditional NLP applications initially used bag-of-words models, which simplify text into isolated words, thus ignoring syntactical structure and sequence Zhang et al. (2010). Though effective in specific contexts like fraud detection Purda and Skillicorn (2015), these methods lack contextual understanding. More recent developments have introduced deep learning-based NLP algorithms, such as Embedding Language Models (ELMo) Mikolov et al. (2013) and Long Short-Term Memory Networks (LSTM) Hochreiter and Schmidhuber (1997), which provide a more effective capture of textual context. Nevertheless, these supervised learning-based methods are highly task-specific and have limited adaptability to generalization Singh et al. (2016).
The emergence of Large Language Models (LLMs) represents a paradigm shift in overcoming these limitations Brown et al. (2020); Touvron et al. (2023); Achiam et al. (2023). Equipped with expansive knowledge bases and sophisticated zero-shot learning capabilities, LLMs are capable of performing a diverse array of text-related tasks—ranging from summarization Zhang et al. (2024) and question-answering Wei et al. (2022) to sentiment analysis Zhang et al. (2023) and E-Commerce Jia et al. (2023) without the need for specialized training in specific tasks or domains Jia et al. (2022); Wang et al. (2023b); Yang et al. (2023); Gruver et al. (2024).
Drawing inspiration from recent advancements in LLMs, this work aims to explore the potential of LLMs to extract trading signals from earnings conference calls (ECCs) to improve predictions of stock performance. ECCs involve senior executives discussing quarterly results, providing a fertile ground for predicting stock movements through nuanced analysis of transcripts and audio. Despite the potential, existing research often overlooks finer details, lacks interpretability, or relies too heavily on sentiment analysis, leading to incomplete data interpretations.
Recent studies have increasingly leveraged domain-specific Large Language Models like FinBERT Liu et al. (2021), which is adapted from Google’s BERT algorithm for financial contexts. FinBERT evaluates the sentiment of ECCs by averaging the sentiment scores of each sentence, revealing a correlation between the sentiment and market reactions. Moreover, the integration of multimodal techniques has significantly improved the accuracy of financial risk predictions. For example, Qin and Yang (2019) and Yang et al. (2020) utilize both textual and auditory data to generate embeddings that encapsulate semantic and auditory features. These models have proven effective in predicting market volatility at various intervals, demonstrating the potential of combining textual analysis with auditory data processing. While these studies may get the descent performance of prediction, another concern with using deep learning-based or LLM techniques is the challenge of explaining how the model arrives at its results. Several interpretability models have been proposed to explain the model’s decision reasons.  Wang et al. (2023a); Tenney et al. (2020).
Existing research, however, reveals significant gaps that warrant further exploration: 1) the previous studies directly input entire texts or audio files into models, potentially missing important details and lacking interpretability, especially regarding which earnings call topics influence predictions; 2) While some studies focus on sentiment extraction from ECCs, they capture only a fraction of the available information, leading to potentially incomplete interpretations. The findings suggest that sentiment analysis alone offers limited explanatory power for predicting stock movements, highlighting the need for broader data utilization; 3) Additionally, integrating large language models (LLMs) into financial analysis while ensuring that investors understand the reasoning behind model outputs poses a distinct challenge, prompting ongoing research.
Given these gaps, in designing a framework to extract trading signals from ECCs to predict financial performance, this paper is interested in exploring the following research questions (RQs):
  • RQ1: How can large language models be used to provide investors with a more comprehensive understanding of a company’s financial health and strategic direction?
  • RQ2: Can a more comprehensive analysis provide additional predictive capability for stock performance?
  • RQ3: Can LLMs be employed to generate interpretable content that aids investors in understanding the decision-making process?
To address the above RQs, this paper introduces the ECC Analyzer, a novel framework utilizing LLMs for in-depth analysis of ECC data. The framework initially provides an understanding of ECCs by segmenting transcripts into themes like financial performance and corporation project discussions. It summarizes these segments, distilling the essence of each thematic chunk, and then combines these summaries into a comprehensive overview. This hierarchical summarization enables stakeholders to grasp complex documents’ main themes and insights. Regarding audio, the model analyzes speech features such as tone, pitch, and intensity to gauge the speaker’s confidence level.
Furthermore, the ECC Analyzer simulates how investors examine key indicators and infer future market behavior. We begin by creating a database (focus) of key indicators with finance experts, such as financial metrics, management changes, operational costs, and strategic plans. Using the constructed database, the ECC Analyzer employs retrieval-augmented generation (RAG) to systematically examine ECCs and pinpoint factors critical to investment decisions. After thoroughly extracting and analyzing the ECC, the results are integrated with raw ECC data to conduct multi-task learning: predicting volatility as well as Value at Risk (VaR) and return for different time intervals (3, 7, 15, and 30 days).
By utilizing RAG, our method improves model interpretability by linking specific earnings call topics directly to stock performance, enhancing both the investor’s understanding and the explainability of the analysis. To validate our approach, we benchmarked our model against traditional methods, showing significant improvements in accuracy and predictive power. These results highlight our model’s utility for financial analysts and investors in making informed decisions based on detailed data analysis.
2 Related Work
This paper integrates concepts from stock performance prediction and advancements in large language models (LLMs). In this section, we offer a concise review of key research in both areas relevant to our study.
2.1 Stock Performance Prediction
The use of NLP techniques to analyze unstructured data for predicting stock performance has attracted significant academic attention. A foundational study by Kogan et al. (2009) shows that simple bag-of-words features from annual reports when combined with historical volatility, can outperform models based solely on historical data. Subsequent research, such as that by Wang and Hua (2014); Rekabsaz et al. (2017); Theil et al. (2018), proposed various document representation methods to predict stock price volatility. Drawing on multimodal technologies,  Qin and Yang (2019) explored how audio features—such as tone, emotion, and speech rate—enhance stock movement predictions when combined with text analysis. Following by this, Yang et al. (2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance. However, the aforementioned studies primarily input ECC data directly into models for prediction without conducting a thorough analysis of the ECC content.
2.2 Large Language Models in Finance
Numerous studies have explored the applications of LLMs in the financial sector.  Li et al. (2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, and 10-Q filings. Yang et al. (2023); Yu et al. (2023) explores the usage of LLMs in mining media news for trading recommendations, showcasing the models’ ability to discern subtle market indicators and sentiments. In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants Lakhani (2023); Subagja et al. (2023); Soni (2023)Abdaljalil and Bouamor (2021); Zmandar et al. (2021) explore the nuanced task of extracting financial and legal items from lengthy text documents, such as financial regulations and comprehensive policy manuals. However, these existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for comprehensive stock performance prediction.
Figure 1: This figure illustrates the ECCs Analyzer’s Framework. The model accepts multimodal inputs: ECC Audio and Transcript. The second area visualizes the model’s pipeline to encode text and audio sources and illustrates how LLMs are applied for information analysis. This model analyzes text-audio pairs, summarizes ECCs, and extracts focal points of interest from a financial expert’s perspective. The third area describes how the model consolidates outputs from both embeddings and LLM analysis for use in subsequent stages. The model will perform multi-task learning: our ECCs Analyzer will predict the Volatility of different terms and VaR in the meantime.
3 Methodology
ECC Analyzer, illustrated in Figure (1) aims to comprehensively understand the multi-data types present in earnings conference calls, including both text and audio components. (3.1) audio encoding (3.2) transcript encoding. Furthermore, it focuses on extracting trading signals from the analyzed data to predict stock performance (3.3) earnings conference call focuses extraction. The model includes a component to optimally integrate different data sources (3.4) additive multi-model fusion. The model performs multi-task prediction: to predict the the volatility across different terms and the value at risk (VaR) at the same time (3.5). It also demonstrates strong predictive ability in forecasting stock returns.
Figure 2: This figure visualizes the mechanism of the earnings conference call analyzer. This framework takes both earnings conference calls’ audio and transcript as input. It starts by hierarchically summarizing the content of ECCs. Further, this model simulates the process through which an investor analyzes an ECC. It examines and extracts focuses on interest, such as financial indicators and business events, from a database designed by financial experts. Subsequently, this model calculates the sentiment for these focal points and extracts audio features, such as changes in tone and pitch, from the corresponding audio segments. All these analyses are combined to form a comprehensive encoder for further use.
3.1 Audio Encoding
Audio pre-trained models have achieved performing results in various downstream tasks Pons and Serra (2019); Cramer et al. (2019); Koh and Dubnov (2021); Wang et al. (2021). We aim to leverage advanced audio pre-trained models like Wav2vec2, a transformer-based Large Language Model recognized for its effectiveness in processing raw audio Baevski et al. (2020), to extract audio embeddings. After that, we employ a Multi-Head Self-Attention (MHSA) mechanism to distill specific audio features. This method is vital for integrating these features with other data modalities, facilitating a more detailed and comprehensive analysis.
To describe the Audio Encoding in more detail, we let the raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample. Each audio frame will be converted into a vector representation: $e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i})$. Therefore, we obtain the audio embeddings $E_{ac}=\{e_{ac}^{1},e_{ac}^{2},\ldots,e_{ac}^{n}\}$ which have dimensions of 520 $\times$ 512, representing the maximum number of audio files across companies and the transform dimensions for a single audio frame, respectively. Audio files with fewer than 520 frames ($n<520$) are zero-padded for consistent matrix size.
$E_{ac}$ are then processed through a MHSA to distill specific audio features. The MHSA includes a multi-head attention block, a norm block, and a two-layer feed-forward network with ReLU activation, forming the basis for all subsequent architectures discussed. In detail, the MHSA calculation process is as follows:
$\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$ (1)
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (2)
where $Q$ (queries) and $K$ (keys) of dimension $d_{k}$ and $V$ values of dimension $d_{v}$. The weights dimensions are: $W_{i}^{Q},W_{i}^{K},W_{i}^{V}\in R^{d_{model}\times d_{k},d_{k},d_{v}}$ respectively, and $W^{o}\in R^{d_{v}\times d_{model}}$. The dot product is then calculated for the query with all the keys. The attention scores are normalized using the softmax function:
$\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$ (3)
The attention function on a set of queries is calculated simultaneously packed together in a matrix Q. The keys and values are also packed in the matrices K and V respectively. Combining (2)-(4), this results in a matrix:
$T_{ac}=\text{MHSA}(E_{ac})$ (4)
where $T_{ac}=\{t_{ac}^{1},t_{ac}^{2},\ldots,t_{ac}^{n}\}$ with size 520 $\times$ 512. $T_{ac}$ is then subjected to an average pooling layer to produce $T_{a}$, a condensed audio feature vector of size 512.
3.2 Transcript Encoding
The transcription process mirrors Audio Encoding, using SimCSE Gao et al. (2021)to extract sentence-level vector representations from earnings conference transcripts. SimCSE is a Siamese neural network architecture that learns to embed pairs of sentences into a shared space where similar sentences are mapped close together and dissimilar sentences are mapped far apart. The raw transcripts are represented as $T_{c}\ =\{t_{c}^{1},a_{c}^{2},\ldots,t_{c}^{n}\}$, with each sentence $t_{c}^{i}$ represents the $i^{th}$ transformed into a vector representation: $e_{tc}^{i}=\text{SimCSE}(t_{c}^{i}).$
We obtain the corresponding text embeddings given by $E_{tc}=\{e_{tc}^{1},e_{tc}^{2},\ldots,e_{tc}^{n}\}$ with size 520 $\times$ 768, where 520 is the maximum number of sentences amongst all data samples and 768 is the dimension of the output of SimCSE. Earnings conference calls with less than 520 sentences ($n<520$) have been zero-padded for uniformity in input matrix size. Same with (1)-(4), the MHSA is applied to $E_{tc}$ to get $T_{tc}=\{t_{tc}^{1},t_{tc}^{2},\ldots,t_{tc}^{n}\}$ with dimension 520 $\times$ 768. Then,$T_{tc}$ is subjected to the average pooling layer to produce $T_{t}$, where $T_{t}$ denotes the resultant extracted textual feature of size 768. This streamlined process effectively captures the textual nuances required for in-depth analysis.
3.3 Earnings Conference Call Focuses Extraction And Analysis
To obtain deep insights from an Earnings Conference Call on how it might influence future market performance, our approach encompasses several steps: (1) summarize ECCs, (2) extract investors’ focus information (3) calculate focus sentiment and (4) extract audio’s features of corresponding focus. Refer to the Figure 2 for additional details.
(1) Summarize ECCs. To accommodate the token limitations of LLMs, we start by dividing the entire document into several chunks and summarizing each one. We then summary these summaries to create a comprehensive overview of the entire document. This two-tiered approach ensures that the summary capture both the detailed and overall information. In further, we will use LLM to get the embedding with a size 1024: $T_{s}=\text{LLM}(summary+chunk\ summaries)$
(2) Extract the investor’s focus using RAG. In this step, we attempt to use LLM to examine and extract focuses of interest to investors. Therefore, we start by identifying the core topics that analysts typically focus on. We consult with four finance experts, each possessing extensive experience in financial analysis. These experts are tasked with meticulously reviewing earnings conference calls across various sectors to pinpoint and summarize the key topics that are frequently discussed and also hold the greatest interest for investors. Their analysis aims to identify overlaps where common discussion topics align closely with investor concerns, highlighting areas of particular significance and interest in the investment community. Once we have established this focused database, our goal is to locate where these topics are mentioned during the earnings calls and systematically extract these segments. However, accurately extracting this information from the dense narrative of earnings conference calls can pose a challenge. Because LLM searches are based on similarity, comparing a single-word topic to an entire document presents a scale issue, and a single word may lack the necessary context, thus diminishing the extraction capability. To enhance the precision of information extraction, we formulate multiple questions related to each topic but phrased differently, ensuring a broad coverage and increasing the likelihood of accurately pinpointing relevant sections. For instance, regarding the dividends topic, we pose questions such as: (1) "Have there been any changes in the stock dividends, and if so, at what rate have the dividends increased or decreased?" (2) Does the company Board expect to increase the stock dividends in the next future? (3) How does the dividend yield compare to the peers?
After locating the segments where these topics are discussed, we extract the corresponding paragraphs. These paragraphs, however, may still contain irrelevant text such as preceding discussions. To address this, we use a contextual compression method from LLM, adept at distilling the paragraph down to its most relevant sentences, effectively obtaining the crucial data from the irrelevant text.
(3) Calculate focus sentiment. Sentiment is a strong indicator for reflecting market perception. We determined to calculate the sentiment score based on the focus. Because this offers a more targeted and insightful view, which may be overlooked in a broader context. It provides a clearer picture of how specific developments or concerns are influencing investor behavior and market movements.
(4) Extract audio’s features using RAG. Once we identify the focal points of interest, we aim to enrich our analysis by locating and examining the audio segments that correspond to these specific topics. By analyzing the audio features related to a particular focus, such as the company’s current projects, we can delve deeper into the emotional nuances and confidence levels exhibited by the speakers. This approach allows us to capture subtle cues in tone, pace, and emphasis that might indicate underlying sentiments or confidence about the discussed topics. The goal is to provide a more layered understanding of how speakers convey their messages and the potential impacts these emotional expressions have on the listeners’ perceptions and reactions. We utilize Praat Boersma and Van Heuven (2001) to extract vocal features, such as pitch, intensity, jitter, HNR(Harmonic to Noise Ratio) and etc, from ECC audio files.
After obtaining the four key components, we then explore effective methods to merge these elements into a cohesive analysis. In further, we will use LLM to get the embedding with size 1024: $T_{f}=\text{LLM}(focus\ analysis+focus\ sentiment+focus\ audio\ features)$
3.4 Additive Multi-modal Fusion
Given the model’s reliance on several inputs and diverse data types, we identify an effective fusion structure to integrate these features into the training process to ensure a balanced weighting among components. We use additive interactions to handle the representational fusion of different abstract representations. These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline Liang et al. (2022). Given the audio feature $T_{a}$, textual feature $T_{t}$ from the transcript, and $T_{s},T_{f}$ from ECC analyzed text, additive fusion can be seen as learning a new joint representation:
$E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{s}+w_{4}\cdot T_{f}+\epsilon$ (5)
where $w_{1}\in R^{512\times 512}$, $w_{2}\in R^{768\times 512}$ and $w_{3},w_{4}\in R^{1024\times 512}$ are the weights learned for additive fusion, $w_{0}$ the bias term and $\epsilon$ the error term. $E$ is a vector with 512 as the final feature from the Earning Conference Call Encoder.
3.5 Multi-Task Prediction
We begin our prediction process by aggregating features from various modules into a comprehensive feature representation. This unified feature set is fed into a two-layer neural network designed to perform the regression task. Integrating these diverse inputs into a cohesive output is crucial, as it harnesses the strengths of each module to enhance analysis and prediction accuracy.
Building on insights from previous research in multimodal financial risk prediction, which has demonstrated substantial improvements in prediction performance through multitask learning, we adopt a joint modeling approach. Here, we concurrently model volatility prediction and VaR prediction using a multi-task framework. The multi-task prediction module is comprised of two separate single-layer feedforward networks, each responsible for predicting volatility (vol) and Value at Risk (var) values individually. We train ECC Analyzer by optimizing multitask loss:
$\mathcal{L}$ $=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})$ (6)
$+(1-\mu)\text{max}(q\times(v-\hat{v}),(1-q)(\hat{v}-v))$
multi-task learning allows us to optimize performance by accurately capturing and predicting multiple stock performance factors simultaneously.
Table 1: Performance results on our proposed framework ECC Analyzer from different baseline models.
Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$ Multi-Task
Classical Method 0.713 1.710 0.526 0.330 0.284 / $\otimes$
LSTM 0.746 1.970 0.459 0.320 0.235 / $\otimes$
MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233 / $\otimes$
HAN 0.598 1.426 0.461 0.308 0.198 / $\otimes$
MRDM 0.577 1.371 0.420 0.300 0.217 / $\otimes$
HTML 0.401 0.845 0.349 0.251 0.158 / D
GPT-4-Turbo 2.198 3.187 5.059 7.959 11.824 0.371 D
ECC Analyzer 0.316 0.553 0.306 0.247 0.159 0.049 D
4 Results and Discussions
4.1 Dataset
The dataset utilized in this study is sourced from the publicly available S&P 500 ECC dataset as constructed by Qin and Yang (2019). It includes both audio recordings and corresponding text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S. stock exchanges. The dataset consists of 572 unique instances where the audio recordings were accurately and closely aligned with the text transcripts. Following the setup by Qin and Yang (2019), we partitioned the dataset into a training set and a test set with an 8:2 ratio, organized temporally to ensure that the data in the training set precedes those in the test set. This temporal division is crucial for maintaining the integrity of our predictive model, aligning the training process with the principle of using historical data to predict future risks—thus enhancing the accuracy and reliability of our forecasting approach.
4.2 Baseline
We compare our approach to several important baselines including 1) GARCH-based Classical Methods; 2) LSTM Gers et al. (2000) model; 3) MT-LSTM+ATT Luong et al. (2015) employing an attention-enhanced LSTM as the foundational model; 4) HAN (Glove) uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels for prediction; 5) MRDM Qin and Yang (2019) proposed a multi-modal deep regression approach for volatility prediction tasks; 6) HTML Yang et al. (2020) presented a state-of-the-art method; and 7) GPT-4-turbo-2024-04-09: directly utilize LLM to predict stock performance. We explain each baseline method in detail in the Appendix A.1
4.3 Implementation Details
We use GPT-4 for the ECC Analyzer experiment, utilizing it to analyze ECC data and build the analysis database. Throughout this process, we set the temperature parameter to 0. This ensures that the Large Language Models (LLMs) produce the most predictable responses, which aids in maintaining consistency in our experiments.
For the overall training of the ECC Analyer framework, we developed the code using PyTorch. Each Multi-Head Attention layer in the network comprises 6 layers and 8 individual heads in each layer. The training process utilized batch sizes $b\in\{2,4,8\}$. We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$. The best hyper-parameters were kept consistent across all experiments, with the exception of the trade-off parameter $\mu$ which varied between the two tasks. We list the evaluation metrics in Appendix A.2.
4.4 Overall Results Analysis (RQ1)
Table 1 shows the performance of various methods in predicting stock performance. Notably, the ECC Analyzer framework excels, especially in short-term and medium-term forecasts, with the lowest Mean Squared Error (MSE) values. Its long-term prediction performance is comparable to the state-of-the-art method, HTML. This improvement highlights the benefits of deep analysis using LLMs extracted from ECCs. However, directly applying LLMs to stock performance prediction proves largely ineffective, similar to random guesses. This indicates that LLMs are more effective as tools to enhance investors’ understanding of a company’s financial health rather than direct predictors of financial metrics.
Additionally, the ECC Analyzer demonstrates outstanding performance in Value at Risk (VaR) prediction, confirming the efficacy of our methodology in providing a nuanced and comprehensive approach to financial risk prediction. We also report the results on Returns in Appendix B.
Table 2: Performance results of ablation study. We designed the ablation study as follows: 1) Audio+Text: uses raw audio and text data from ECCs; 2)Audio+Text+$E_{os}$: adds a comprehensive ECC summary generated by LLMs; 3) Audio+Text+$E_{os}+E_{cs}$: integrates both overall and chunk summaries for the ECC; 4) Audio+Text+$E_{fo}$: combines raw data with focused analytical results; 5) $E_{os}+E_{cs}+E_{fo}$: merges all LLM analyses for prediction without raw data; 6) Audio+Text+$E_{os}+E_{cs}+E_{fo}$: combines all data and analyses for enhanced prediction.
Module $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$
Audio+Text 0.373 0.645 0.362 0.28 0.204
Audio+Text+$E_{os}$ 0.373 0.638 0.380 0.276 0.201
Audio+Text+$E_{os}+E_{cs}$ 0.357 0.627 0.335 0.267 0.199
Audio+Text+$E_{fo}$ 0.324 0.579 0.323 0.23 0.165
$E_{os}+E_{cs}+E_{fo}$ 0.343 0.601 0.344 0.247 0.179
Audio+Text+$E_{os}+E_{cs}+E_{fo}$ 0.310 0.553 0.306 0.22 0.159
4.5 Ablation Study (RQ2)
In our research, we conducted an ablation study to assess how different combinations of ECC analysis results impact our model’s performance. This systematic comparison helped us identify the individual contributions of each component.
According to Table 2, we can find that the audio and text features extracted by advanced large language models significantly improved short-term prediction accuracy compared to previous methods. Furthermore, incorporating summaries of the data slightly enhanced performance, but more notable improvements were observed when we added analysis of specific focus points. This indicates that our model effectively isolates and utilizes the most relevant information for predicting stock movements. Our best analytical results come from integrating the full spectrum of data and analytical outputs, underscoring the value of each component in our model. Notably, good predictive results were also can obtained using only analytics derived from LLMs, affirming the response to our RQ2: comprehensive analysis indeed enhances the predictive capability for stock performance.
Our findings also suggest that while earnings calls are information-rich, including every detail in the analysis can be counterproductive and may cloud essential insights. It is therefore critical to pinpoint and concentrate on the most predictive elements of the data, filtering out less relevant information to optimize the analysis process for stock performance prediction.
4.6 Interpretability Study (RQ3)
In our study, we initiate the analysis by collaborating with consultants to summarize and identify key focus points from earnings conference calls; the more comprehensive the summary, the better. Following this initial step, we combine these focus points and try different combinations to determine the optimal subset that most strongly predicts stock performance. (The finalized list of focus points is included in the appendix.) This method serves a dual purpose for investors. Firstly, it reveals which focus points are highly relevant to stock performance, allowing investors to prioritize their attention effectively. Secondly, by extracting and analyzing these key focuses, investors can understand the specific contents in the ECC that drive movements in the stock market.
5 Conclusions
This study confirms the effectiveness of using Large Language Models (LLMs) to extract and analyze key topics from earnings conference calls. Our approach not only pinpoints critical discussions but also assesses their impact on stock performance, predicting metrics like volatility and Value at Risk (VaR). By integrating textual and audio data, our model offers a comprehensive view, capturing subtleties such as tone and pitch. Results from benchmark comparisons demonstrate our model’s superior accuracy and predictive capabilities, highlighting LLMs’ potential to improve interpretability and decision-making in investments.
Appendix A Supplement Material for Experiment Setup
A.1 Baseline Setup
  • Classical Methods: We incorporate the GARCH model and its derivatives, as described in  Franses and Van Dijk (1996); Kim and Won (2018). These models are well-recognized for short-term volatility prediction but may not be as effective for forecasting average volatility over longer periods, such as n-day volatility.
  • LSTM Gers et al. (2000): Long Short-Term Memory Networks (LSTMs) are a popular choice for financial time series prediction due to their efficacy in handling sequential data. We use a straightforward LSTM model as a benchmark for volatility prediction.
  • MT-LSTM+ATT Luong et al. (2015): combines the prediction of average n-day volatility with the forecasting of single-day volatility, employing an attention-enhanced LSTM as the foundational model.
  • HAN (Glove): uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels. HAN first get word embeddings using pre-trained Glove vectors and then processed by a Bi-GRU Chung et al. (2014) encoder, while another Bi-GRU encoder simultaneously forms a sentence-level representation of each document. The resulting document representation is input into a regression layer to produce predictions.
  • MRDM Qin and Yang (2019): The MRDM model first introduced a multi-modal deep regression approach to fuse the GloVe embeddings and hand-crafted acoustic features for volatility prediction tasks.
  • HTML Yang et al. (2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding. Similar to MDRM, HTML also leverages the same audio features. These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
  • GPT-4-turbo-2024-04-09: We assessed the capability of LLMs in directly predicting stock performance from ECCs. The model was set to generate response with a zero temperature setting to ensure deterministic output.
A.2 Evalation Metrics
A.3 Volatility of Different Future Intervals
We used the following formula to calculate the Volatility of Different Terms:
$\sigma_{ijk}^{2}=\frac{1}{k-1}\sum_{m=1}^{k}[S_{i,(j-m)}-\frac{1}{k}\sum_{l=1}^{k}S_{i,(j-l)}]^{2}$
Here, we define the $\sigma_{ijk}$ as the $k-days$ volatility of Stock $i$ at time $j$, which is calculated as the sample standard deviation of the past $k-days$ stock closed price of company $i$. But in the real application, we took the log form of historical volatility.
A.4 VaR of Different Future Intervals
Our second task is predicting the 1-day VaR of the target stock based on the multi-source inputs. The definition of VaR is:
$x=F^{-1}(p)$
The $F(\cdot)$ is the cumulative loss distribution, $p$ is the percentile we set, and $x$ is the VaR. From the idea of Quantile Regression, we can have:
$L_{\tau}(y,\hat{y})=\begin{cases}\tau\cdot(y-\hat{y})&\text{if }y\geq\hat{y}\\ (1-\tau)\cdot(\hat{y}-y)&\text{if }y<\hat{y}\end{cases}$
Calculating and estimating VaR can help the company better deal with financial risks and avoid extreme scenarios in the future.
A.5 Returns with Different Future Intervals
We used the following formula to calculate the Returns with Different Gaps:
$R_{ijk}=\frac{S_{i,(j+k)}-S_{i,j}}{S_{i,j}}$
Here, we denote $S_{i,j}$ as the closed price of company $i$’s stock at time $j$, and $R_{ijK}$ as the $k-day$ return of stock price $i$ at time $j$. We calculated this value as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$.
Appendix B The Results on Returns
To evaluate the predictive capability of our model, we have applied it to forecast stock returns. The results of these forecasts are systematically compiled and displayed in a table 3.
Table 3: ECC Analyzer’s Prediction Performance On Stock Returns
Return_3d Return_7d Return_15d Return_30d
Predited Return 0.0024 0.0068 0.0131 0.0185
True Return 0.0007 0.0017 0.0035 0.0139
Error 0.0018 0.0051 0.0096 0.0046
Percentage Error 2.6923 3.0964 2.7216 0.3348
Appendix C Regression and Sentiment Analysis
In the main chapter, we used the various returns and volatility of different terms as our predicted target. This part will introduce how we calculated these values and further results about the relationships between sentiments calculated by LLM and these target values.
C.1 Cumulative Abnormal Returns and Regression Analysis
We used both Abnormal Returns collected from the market and the returns of different terms to capture the sentiment delivered by ECCs. For the abnormal Returns, we used the following regression model Lee and Connolly (2010):
$R_{it}=\alpha+\beta R_{mt}+\varepsilon_{it}$
$R_{it}$ is the return of the security at time $t$, also $R_{mt}$ is the market portfolio return at time $t$. The Abnormal Returns we defined here are the difference between the real value of Returns and estimated ones, which is $\varepsilon_{it}$. For other dependent variables, we used the $R_{ijk}$ for $k=1,3,7,15,30$. Results are shown in Table 4.
Table 4: This table shows the linear regression results of different kinds of returns. We used different types of returns as dependent variables. $CARs$ is the cumulative abnormal returns, calculated as the difference between the actual value of the predicted valve from factor models. The rest of the dependent variables are calculated as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$. The independent variable is the Sentiment score, which is calculated by LLM. We extracted all the events in one ECC and evaluated their sentiment score. The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
Dependent Variable
CARs Return_1d Return_3d Return_7d Return_15d Return_30d
Constant $-0.0105^{***}$ $-0.0105^{***}$ $-0.0126^{***}$ $-0.0182^{***}$ $-0.0286^{***}$ $-0.0244^{***}$
(0.004) (0.004) (0.004) (0.005) (0.007 (0.008)
Sentiment $0.0223^{***}$ $0.024^{***}$ $0.0295^{***}$ $0.0442^{***}$ $0.0716^{***}$ $0.0852^{***}$
(0.008) (0.008) (0.009) (0.010) (0.015) (0.016)
Adjusted $R^{2}$ 0.016 0.018 0.022 0.038 0.048 0.055
Observation 572 572 572 572 572 572
For the volatility, we used the difference between the two closed periods of volatility. The dependent variable for volatility could be calculated as $log(\sigma_{i(j+1)k})-log(\sigma_{ijk})$ for $k=3,7,15,30$. Results are shown in Table 5.
Table 5: This table shows the linear regression results of different terms of Volatility. We used four different terms of volatility. Volatility here is defined by the log form of the sample standard deviation of the past $k-days$ stock closed price of company $i$. The independent variable is the Sentiment score, which is calculated by LLM. We extracted all the events in one ECC and evaluated their sentiment score. The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
Dependent Variable
Volatility_3d Volatility_7d Volatility_15d Volatility_30d
Constant $-0.0254^{***}$ $-0.017^{***}$ $-0.0071^{***}$ $-0.005^{***}$
(0.013) (0.009) (0.006) (0.004)
Sentiment $-0.0681^{***}$ $-0.0556^{***}$ $-0.0438^{***}$ $-0.0266^{***}$
(0.026) (0.018) (0.013) (0.009)
Adjusted $R^{2}$ 0.013 0.019 0.022 0.019
Observation 572 572 572 572
From Table 4 and Table 5, we can find that the Sentiment generated from the LLM is always statistically significant in the linear regression model results. Comparing the coefficient of different terms of Returns, we found that the coefficient gets larger when there are more significant term gaps. This indicates that the impact of the ECC’s sentiment will appear as time progresses. There is a higher probability that the impact of the ECCs occurs after a certain period because, besides the financial data, most information, like future projects mentioned in ECCs, will not be reflated instantly. That could explain why the coefficient will rise as the period gets larger. However, the volatility change will have a smaller impact when the period gets larger(because the absolute value of Sentiments’ coefficient in Table 5 is smaller when the period gets more extensive), which means that the same ECCs sentiments effect will have a more significant impact in short terms rather than the longer terms.
Figure 3: We asked four experts to help us analyze the possible scenarios of ECCs. Our experts focus on the following general Categories of the ECCs’ transcript: Financial Indicators, Employee & Managers, Cost, Project Expansion, Business, and Future Look. Each general category is separated into some minor focusing points. Our experts put forward three questions for each small focusing point. We used these questions to ask LLM and wished for a directional response in the thinking pattern hidden in the questions that we wanted to teach the LLM.
Appendix D Prompt Design
D.1 Prompt for Summarizing Earnings Conference Call Segments
  • Identify Key Points
    For each segment, identify the key topics covered. Note any significant financial figures, strategic decisions, performance metrics, or forward-looking statements.
  • Summarize Succinctly
    Write a concise summary for each segment, capturing the essence of the discussion. Aim to condense the information into a few sentences that clearly convey the main points and outcomes discussed.
  • Highlight Relevant Details
    Include any specific details that are critical for understanding the segment’s context or implications, such as notable quotes from the company’s executives or specific data points that illustrate trends or changes.
  • Connect the Dots
    If applicable, relate the segment’s content to broader company objectives or industry trends to provide context and show how the segment fits into the bigger picture.
D.2 Prompt for Creating an Overview Summary from Earnings Conference Call Segments
  • Gather Segment Summaries
    Start by reviewing the summaries of each segment from the earnings conference call. Ensure that you have all the segment summaries available to reference.
  • Identify Common Themes
    Look for common themes, recurring issues, or consistent messages across the segments. Note any overarching strategies, goals, or concerns expressed by the company executives.
D.3 Prompt for Extracting Focus And Explore Its Impact
  • Listen to the Call
    Begin by thoroughly listening to the entire earnings conference call. Pay attention to both the prepared remarks and the question-and-answer session.
  • Identify Focus Points
    Identify statements or discussions that involve significant financial metrics, strategic initiatives, new products or markets, regulatory impacts, or any notable shifts in operations. These are potential focus points that could influence investor perceptions and stock price.
  • Document Evidence
    For each identified focus point, document the exact wording used, the context in which it was discussed, and who discussed it (e.g., CEO, CFO). This will be crucial for accurate interpretation and analysis.
  • Analyze Impact on Stock Movement
    Pre and Post Analysis: Examine stock price movements immediately before and after the call to capture initial reactions.
  • Longer-term Impact
    Review stock performance in the days or weeks following the call to assess sustained impacts.
  • Compare with Market Trends
    Ensure to factor in overall market conditions and sector movements to isolate the impact of the earnings call from broader market trends.
D.4 Prompt for Analyzing Sentiment of Focus Points and Supporting Evidence from an Earnings Conference Call
  • Analyze Sentiment for Each Focus Point
    Apply the sentiment analysis to the text surrounding each focus point. Pay attention to the language used, such as positive, negative, or neutral descriptors, and the intensity of the language. Please assign a continuous decimal sentiment score to each event, ranging from -1 to 1. A score of -1 represents a highly negative sentiment, 0 indicates neutrality, and 1 signifies a highly positive sentiment.
  • Context Consideration
    Consider the context in which each point was discussed. Assess whether the sentiment is directly related to the focus point or influenced by broader discussion themes.
  • Document Supporting Quotes
    For each focus point, document specific quotes or statements from the call that illustrate the sentiment. Note the speaker and their role to add credibility to the sentiment analysis.
D.5 Prompt for Analyzing Audio Features of Focus Points from an Earnings Conference Call
  • Tone and Pitch Analysis
    Examine variations in tone and pitch within each audio segment. Look for patterns that might indicate emphasis, uncertainty, confidence, or stress.
  • Volume and Speech Rate
    Measure changes in volume and variations in speech rate. High volume and rapid speech may indicate areas of strong emotion or importance.
  • Pause Patterns
    Identify the frequency and duration of pauses, which can provide insights into the speaker’s thought process or hesitation.

### v2

@@
-ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Performance Prediction
-Abstract
-In the realm of financial analytics, leveraging unstructured data, such as earnings conference calls (ECCs), to forecast stock performance is a critical challenge that has attracted both academics and investors.
-While previous studies have used deep learning-based models to obtain a general view of ECCs, they often fail to capture detailed, complex information.
-Our study introduces a novel framework: ECC Analyzer, combining Large Language Models (LLMs) and multi-modal techniques to extract richer, more predictive insights.
-The model begins by summarizing the transcript’s structure and analyzing the speakers’ mode and confidence level by detecting variations in tone and pitch for audio.
-This analysis helps investors form an overview perception of the ECCs.
-Moreover, this model uses the Retrieval-Augmented Generation (RAG) based methods to meticulously extract the focuses that have a significant impact on stock performance from an expert’s perspective, providing a more targeted analysis.
-The model goes a step further by enriching these extracted focuses with additional layers of analysis, such as sentiment and audio segment features.
-By integrating these insights, the ECC Analyzer performs multi-task predictions of stock performance, including volatility, value-at-risk (VaR), and return for different intervals.
-The results show that our model outperforms traditional analytic benchmarks, confirming the effectiveness of using advanced LLM techniques in financial analytics.
-1 Introduction
-The integration of structured data such as stock prices and financial ratios with unstructured data including financial filings and company news is increasingly essential in investment decision-making Fang and Zhang (2016); Roeder et al.
-(2022).
-This trend stems from the recognition that unstructured data can provide insights not fully captured by structured data alone.
-For instance, financial reports elucidate the broader implications of numerical data through discussions of managerial decisions and corporate strategies Wang and Hua (2014), while company news can shed light on public sentiment, emerging market trends, and market perceptions Kogan et al.
-(2009); Tetlock (2007).
-These nuanced insights are vital for a comprehensive analysis of complex market dynamics.
-Advancements in Natural Language Processing (NLP) have significantly enhanced the ability to analyze unstructured data within the financial sector.
-Traditional NLP applications initially used bag-of-words models, which simplify text into isolated words, thus ignoring syntactical structure and sequence Zhang et al.
-(2010).
-Though effective in specific contexts like fraud detection Purda and Skillicorn (2015), these methods lack contextual understanding.
-More recent developments have introduced deep learning-based NLP algorithms, such as Embedding Language Models (ELMo) Mikolov et al.
-(2013) and Long Short-Term Memory Networks (LSTM) Hochreiter and Schmidhuber (1997), which provide a more effective capture of textual context.
-Nevertheless, these supervised learning-based methods are highly task-specific and have limited adaptability to generalization Singh et al.
-(2016).
-The emergence of Large Language Models (LLMs) represents a paradigm shift in overcoming these limitations Brown et al.
-(2020); Touvron et al.
-(2023); Achiam et al.
-(2023).
-Equipped with expansive knowledge bases and sophisticated zero-shot learning capabilities, LLMs are capable of performing a diverse array of text-related tasks—ranging from summarization Zhang et al.
-(2024) and question-answering Wei et al.
-(2022) to sentiment analysis Zhang et al.
-(2023) and E-Commerce Jia et al.
-(2023) without the need for specialized training in specific tasks or domains Jia et al.
-(2022); Wang et al.
-(2023b); Yang et al.
-(2023); Gruver et al.
-(2024).
-Drawing inspiration from recent advancements in LLMs, this work aims to explore the potential of LLMs to extract trading signals from earnings conference calls (ECCs) to improve predictions of stock performance.
-ECCs involve senior executives discussing quarterly results, providing a fertile ground for predicting stock movements through nuanced analysis of transcripts and audio.
-Despite the potential, existing research often overlooks finer details, lacks interpretability, or relies too heavily on sentiment analysis, leading to incomplete data interpretations.
-Recent studies have increasingly leveraged domain-specific Large Language Models like FinBERT Liu et al.
-(2021), which is adapted from Google’s BERT algorithm for financial contexts.
-FinBERT evaluates the sentiment of ECCs by averaging the sentiment scores of each sentence, revealing a correlation between the sentiment and market reactions.
-Moreover, the integration of multimodal techniques has significantly improved the accuracy of financial risk predictions.
-For example, Qin and Yang (2019) and Yang et al.
-(2020) utilize both textual and auditory data to generate embeddings that encapsulate semantic and auditory features.
-These models have proven effective in predicting market volatility at various intervals, demonstrating the potential of combining textual analysis with auditory data processing.
-While these studies may get the descent performance of prediction, another concern with using deep learning-based or LLM techniques is the challenge of explaining how the model arrives at its results.
-Several interpretability models have been proposed to explain the model’s decision reasons.
-Wang et al.
-(2023a); Tenney et al.
-(2020).
-Existing research, however, reveals significant gaps that warrant further exploration: 1) the previous studies directly input entire texts or audio files into models, potentially missing important details and lacking interpretability, especially regarding which earnings call topics influence predictions; 2) While some studies focus on sentiment extraction from ECCs, they capture only a fraction of the available information, leading to potentially incomplete interpretations.
-The findings suggest that sentiment analysis alone offers limited explanatory power for predicting stock movements, highlighting the need for broader data utilization; 3) Additionally, integrating large language models (LLMs) into financial analysis while ensuring that investors understand the reasoning behind model outputs poses a distinct challenge, prompting ongoing research.
-Given these gaps, in designing a framework to extract trading signals from ECCs to predict financial performance, this paper is interested in exploring the following research questions (RQs):
-  • RQ1: How can large language models be used to provide investors with a more comprehensive understanding of a company’s financial health and strategic direction?
-  • RQ2: Can a more comprehensive analysis provide additional predictive capability for stock performance?
-  • RQ3: Can LLMs be employed to generate interpretable content that aids investors in understanding the decision-making process?
-To address the above RQs, this paper introduces the ECC Analyzer, a novel framework utilizing LLMs for in-depth analysis of ECC data.
-The framework initially provides an understanding of ECCs by segmenting transcripts into themes like financial performance and corporation project discussions.
-It summarizes these segments, distilling the essence of each thematic chunk, and then combines these summaries into a comprehensive overview.
-This hierarchical summarization enables stakeholders to grasp complex documents’ main themes and insights.
-Regarding audio, the model analyzes speech features such as tone, pitch, and intensity to gauge the speaker’s confidence level.
-Furthermore, the ECC Analyzer simulates how investors examine key indicators and infer future market behavior.
-We begin by creating a database (focus) of key indicators with finance experts, such as financial metrics, management changes, operational costs, and strategic plans.
-Using the constructed database, the ECC Analyzer employs retrieval-augmented generation (RAG) to systematically examine ECCs and pinpoint factors critical to investment decisions.
-After thoroughly extracting and analyzing the ECC, the results are integrated with raw ECC data to conduct multi-task learning: predicting volatility as well as Value at Risk (VaR) and return for different time intervals (3, 7, 15, and 30 days).
-By utilizing RAG, our method improves model interpretability by linking specific earnings call topics directly to stock performance, enhancing both the investor’s understanding and the explainability of the analysis.
-To validate our approach, we benchmarked our model against traditional methods, showing significant improvements in accuracy and predictive power.
-These results highlight our model’s utility for financial analysts and investors in making informed decisions based on detailed data analysis.
-2 Related Work
-This paper integrates concepts from stock performance prediction and advancements in large language models (LLMs).
-In this section, we offer a concise review of key research in both areas relevant to our study.
-2.1 Stock Performance Prediction
-The use of NLP techniques to analyze unstructured data for predicting stock performance has attracted significant academic attention.
-A foundational study by Kogan et al.
-(2009) shows that simple bag-of-words features from annual reports when combined with historical volatility, can outperform models based solely on historical data.
-Subsequent research, such as that by Wang and Hua (2014); Rekabsaz et al.
-(2017); Theil et al.
-(2018), proposed various document representation methods to predict stock price volatility.
-Drawing on multimodal technologies,  Qin and Yang (2019) explored how audio features—such as tone, emotion, and speech rate—enhance stock movement predictions when combined with text analysis.
-Following by this, Yang et al.
-(2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance.
+ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Volatility Prediction
+Abstract.
+In the realm of financial analytics, leveraging unstructured data, such as earnings conference calls (ECCs), to forecast stock volatility is a critical challenge that has attracted both academics and investors.
+While previous studies have used multimodal deep learning-based models to obtain a general view of ECCs for volatility predicting, they often fail to capture detailed, complex information.
+Our research introduces a novel framework: ECC Analyzer, which utilizes large language models (LLMs) to extract richer, more predictive content from ECCs to aid the model’s prediction performance.
+We use the pre-trained large models to extract textual and audio features from ECCs and implement a hierarchical information extraction strategy to extract more fine-grained information.
+This strategy first extracts paragraph-level general information by summarizing the text and then extracts fine-grained focus sentences using Retrieval-Augmented Generation (RAG).
+These features are then fused through multimodal feature fusion to perform volatility prediction.
+Experimental results demonstrate that our model outperforms traditional analytical benchmarks, confirming the effectiveness of advanced LLM techniques in financial analysis.
+Large Language Model, Earnings Conference Call Analysis, Volatility forecasting, Retrieval-Augmented Generation
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Information systems Multimedia information systems
+1. Introduction
+Predicting the stock volatility over a certain period is a crucial task in financial analysis, aiding capital market participants in making better investment decisions.
+As such, developing effective techniques for predicting stock volatility has become increasingly important among academics and the financial industry.
+Previous studies in economics have shown that stock volatility can be predicted using publicly available information (Dumas et al., 2009).
+Substantial research in finance and computational linguistics has made significant progress in predicting stock volatility from various textual sources, such as company disclosed reports (Kogan et al., 2009; Hoberg and Phillips, 2016; Rekabsaz et al., 2017), the transcripts of earnings conference calls (Kimbrough, 2005; Wang and Hua, 2014) and social media (Bollen et al., 2011; Ding et al., 2015; Oliveira et al., 2017).
+Recently, advancements in multimodal learning have enabled the use of more unstructured multimedia data, such as audio recordings of earning conference calls (Qin and Yang, 2019; Li et al., 2020), Merge & Acquisition calls (Sawhney et al., 2021) and the video of CEO’s speech (Rajandran, 2021), for stock volatility prediction.
+Multimodal learning is particularly valuable in this context as it allows for the integration of diverse data sources, providing richer, more nuanced insights into market dynamics and sentiment.
+This comprehensive analysis is essential for understanding complex market behaviors (Qin and Yang, 2019).
+Our study focuses on multimodal earning conference calls (ECCs) data for two primary reasons: 1) transcripts and audio recordings are publicly available, and 2) ECCs are often associated with high volatility primarily due to the market reaction to the earnings announcement (Foster et al., 1984).
+Existing multimodal approaches to volatility prediction using ECCs typically extract features from speech and text separately.
+Subsequently, commonly used Natural Language Processing (NLP) models, such as LSTM or Transformer-based models, are employed to jointly model the extracted speech and text features, ultimately performing volatility prediction (Qin and Yang, 2019; Yang et al., 2020; Wang et al., 2024).
+While these methods have shown that the multimodal approach can extract complementary information from multiple modalities, enhancing financial modeling performance and model robustness, several challenges remain unaddressed: firstly, previous work directly feeds features extracted from text and audio into the model, potentially missing important contextual details; secondly, these approaches assign equal importance to all sentences and audio segments which hard to reflect the impact of important information in the ECCs on the prediction.
+Figure 1.
+illustrates the ECC Analyzer Framework.
+The proposed method accepts multimodal inputs: audio record and transcript.
+The upper part of the box illustrates the feature extraction process for both the audio and text of the data.
+We use the pre-trained large models to generate embeddings from the audio and text, followed by a transformer encoder to extract the corresponding features.
+The lower part of the box represents a deeper analysis of the ECC.
+Here, the text is divided into chunks, which are then summarized into paragraphs using an LLM.
+Key sentences are extracted via RAG and converted into text features through text embedding.
+These text features are then fused with the features extracted from the upper part of the box to make the final volatility prediction.
+In response to the aforementioned limitations and inspired by the significant performance improvements offered by large language models (LLMs) in various downstream NLP tasks, this paper introduces the ECC Analyzer — a novel framework that leverages LLMs for in-depth analysis of ECC data to enhance volatility prediction performance.
+The proposed ECC Analyzer employs a hierarchical information extraction strategy: 1) advanced pre-trained large models are used to extract embeddings from both audio and text, capturing the overall information content of ECCs.; 2) LLMs summarize ECC transcripts at the paragraph level, distilling important information from the transcript; 3) In collaboration with financial experts, we designed the “Question Bank” which contains a series of questions about ECC content that are of interest to investors.
+These questions are used as queries to extract fine-grained sentences containing key information from ECCs using Retrieval-Augmented Generation (RAG).
+We use these questions as queries and extract fine-grained sentences containing key information from ECC by RAG to improve the prediction performance of the model.
+Then the summarized text with extracted key focus sentences is transformed into text embedding and combined with the extracted audio and text features for volatility prediction.
+Our extensive experiments on the real-world S&P 500 ECCs dataset demonstrate clear and significant improvements in prediction accuracy attributable to our proposed approach.
+We claim that the contribution is two-fold: 1) Our framework innovatively integrates multi-modal information with fine-grained details extracted by Large Language Models (LLMs) to effectively distill and utilize rich information from financial documents, such as ECCs, for prediction tasks.
+2) Our pipeline demonstrates a $27.7\%$ reduction in average Mean Squared Error (MSE) compared to the current state-of-the-art (SOTA) model.
+In detail, The results show that our method achieves substantial performance improvements in forecasting short-term volatility (3-day, 7-day) compared to current SOTA methods.
+The results for medium-term volatility (15 days) and long-term volatility (30 days) are comparable to the best existing methods.
+2. Related Work
+2.1. Stock Volatility Prediction
+Volatility forecasting has long been of interest to researchers due to its practical applications.
+Conventional forecasting approaches rely on historical stock prices and use continuous and discrete time series models (Kristjanpoller et al., 2014; Cox and Ross, 1976; Heston, 1993; Hsu et al., 2008; Franses and Van Dijk, 1996; Kim and Won, 2018).
+Recently, the use of NLP techniques to analyze unstructured data for predicting stock performance has attracted significant academic attention.
+A foundational study by (Kogan et al., 2009) shows that simple bag-of-words features from annual reports when combined with historical volatility, can outperform models based solely on historical data.
+Subsequent research, such as that by (Wang and Hua, 2014; Rekabsaz et al., 2017; Theil et al., 2018), proposed various document representation methods to predict stock price volatility.
+Drawing on multimodal technologies,  (Qin and Yang, 2019) explored how audio features—such as tone, emotion, and speech rate—enhance stock movement predictions when combined with text analysis.
+Following by this, (Yang et al., 2020) further extends the idea of using multimodal data to improve risk prediction performance in multi-task learning, and the authors’ experiments show that predicting multiple tasks at the same time can help the model further improve prediction performance.
+(Wang et al., 2024) addressed the reduction of gender bias in ECC predictions through adversarial training.
 However, the aforementioned studies primarily input ECC data directly into models for prediction without conducting a thorough analysis of the ECC content.
-2.2 Large Language Models in Finance
+2.2. Large Language Models in Financial Application
 Numerous studies have explored the applications of LLMs in the financial sector.
-Li et al.
-(2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, and 10-Q filings.
-Yang et al.
-(2023); Yu et al.
-(2023) explores the usage of LLMs in mining media news for trading recommendations, showcasing the models’ ability to discern subtle market indicators and sentiments.
-In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants Lakhani (2023); Subagja et al.
-(2023); Soni (2023)Abdaljalil and Bouamor (2021); Zmandar et al.
-(2021) explore the nuanced task of extracting financial and legal items from lengthy text documents, such as financial regulations and comprehensive policy manuals.
-However, these existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for comprehensive stock performance prediction.
-Figure 1: This figure illustrates the ECCs Analyzer’s Framework.
-The model accepts multimodal inputs: ECC Audio and Transcript.
-The second area visualizes the model’s pipeline to encode text and audio sources and illustrates how LLMs are applied for information analysis.
-This model analyzes text-audio pairs, summarizes ECCs, and extracts focal points of interest from a financial expert’s perspective.
-The third area describes how the model consolidates outputs from both embeddings and LLM analysis for use in subsequent stages.
-The model will perform multi-task learning: our ECCs Analyzer will predict the Volatility of different terms and VaR in the meantime.
-3 Methodology
-ECC Analyzer, illustrated in Figure (1) aims to comprehensively understand the multi-data types present in earnings conference calls, including both text and audio components.
-(3.1) audio encoding (3.2) transcript encoding.
-Furthermore, it focuses on extracting trading signals from the analyzed data to predict stock performance (3.3) earnings conference call focuses extraction.
-The model includes a component to optimally integrate different data sources (3.4) additive multi-model fusion.
-The model performs multi-task prediction: to predict the the volatility across different terms and the value at risk (VaR) at the same time (3.5).
-It also demonstrates strong predictive ability in forecasting stock returns.
-Figure 2: This figure visualizes the mechanism of the earnings conference call analyzer.
-This framework takes both earnings conference calls’ audio and transcript as input.
-It starts by hierarchically summarizing the content of ECCs.
-Further, this model simulates the process through which an investor analyzes an ECC.
-It examines and extracts focuses on interest, such as financial indicators and business events, from a database designed by financial experts.
-Subsequently, this model calculates the sentiment for these focal points and extracts audio features, such as changes in tone and pitch, from the corresponding audio segments.
-All these analyses are combined to form a comprehensive encoder for further use.
-3.1 Audio Encoding
-Audio pre-trained models have achieved performing results in various downstream tasks Pons and Serra (2019); Cramer et al.
-(2019); Koh and Dubnov (2021); Wang et al.
-(2021).
-We aim to leverage advanced audio pre-trained models like Wav2vec2, a transformer-based Large Language Model recognized for its effectiveness in processing raw audio Baevski et al.
-(2020), to extract audio embeddings.
+(Li et al., 2023) explore how LLMs have been adeptly applied to summarize and abstract complex financial documents such as 10-K, and 10-Q filings.
+The FinBen (Xie et al., 2024) provides the benchmark performance of LLM for each task in the financial domain.
+FinGPT (Yang et al., 2023) and FinMem (Yu et al., 2023) explore the usage of LLMs in mining media news for trading recommendations, showcasing the models’ ability to discern subtle market indicators and sentiments.
+In the domain of customer service, the implementation of LLM-powered chatbots is spotlighted for offering context-aware interactions, serving as both assistants and consultants (Lakhani, 2023; Subagja et al., 2023; Soni, 2023)(Abdaljalil and Bouamor, 2021; Zmandar et al., 2021) explore the nuanced task of extracting financial and legal items from lengthy text documents, such as financial regulations and comprehensive policy manuals.
+However, these existing studies predominantly focus on tasks like financial text summarization, question-answering (Q&A), and stock movement prediction (binary classification), with a notable gap in the application of LLMs for comprehensive stock volatility prediction.
+3. Problem Formulation
+Volatility can be expressed as the natural log of the standard deviation of return prices $r$ in a window of ${\tau}$ days.
+We calculate the 3, 7, 15, and 30 trading days volatility using the equation:
+(1) $v_{[d-\tau,d]}=\ln\left(\left(\frac{\sum_{i=0}^{\tau}(r_{d-i}-\bar{r})^{2}}{\tau}\right)^{\frac{1}{2}}\right)$
+The notations used across the paper are discussed herein.
+Let $s\in S$ denote a stock, $c\in C$ be an earnings call for stock s.
+For each stock, there exist multiple earnings calls c that are held periodically.
+Each call c can be segmented into a set of $a_{c}^{i}\in A_{c}$ audio clips, and corresponding $t_{c}^{i}\in T_{c}$ text sentences for $i\in[1,N]$, where is the $N$ is the maximum number of audio clips in a call.
+For each stock, there exists a daily return denoted by $r_{i}=\ \frac{p_{i}-p_{i-1}}{p_{i-1}}$, where $p_{i}$ is the adjusted close price at the end of the day and r̄ is the average return over a period of ${\tau}$ days.
+We aims to develop a predictive regression function $f(c)\rightarrow v_{[d-\tau,d]}$
+For the evaluation, following (Qin and Yang, 2019), we assess the accuracy of volatility predictions by comparing the predicted values $y_{i}$ with the labeled volatility values $\hat{y_{i}}$.
+We use Mean Squared Error (MSE) as the evaluation metric:
+(2) $MSE=\frac{\sum_{i}(y_{i}-\hat{y_{i}})^{2}}{n}$
+This metric quantifies the average squared difference between the predicted and actual volatility values, providing a measure of prediction accuracy.
+Additionally, for constructing volatility labels, we collected stock price information via the Yahoo Finance API [^1] https://finance.yahoo.com/ and used equation 1 to calculate the labels for days 3, 7, 15, and 30.
+4. Our proposed framework
+ECC Analyzer (in Figure 1) aims to comprehensively understand the multi-data types present in earnings conference calls, including both text and audio components.
+In this section, we explain our methodology in 4 parts: (4.1) audio encoding, (4.2) transcript encoding,(4.3) fine-grained information extraction from the transcript by using LLM, and (4.4) multi-model fusion and model training.
+4.1. Audio Encoding
+Audio pre-trained models have achieved performing results in various downstream tasks (Pons and Serra, 2019; Cramer et al., 2019; Koh and Dubnov, 2021; Wang et al., 2021).
+We aim to leverage advanced audio pre-trained models like Wav2vec2, a transformer-based Large Language Model recognized for its effectiveness in processing raw audio (Baevski et al., 2020), to extract audio embeddings.
 After that, we employ a Multi-Head Self-Attention (MHSA) mechanism to distill specific audio features.
 This method is vital for integrating these features with other data modalities, facilitating a more detailed and comprehensive analysis.
-To describe the Audio Encoding in more detail, we let the raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample.
+Figure 2. visualizes the process of fine-grained information extraction from ECC transcript.
+The raw audio input data be represented by $A_{c}=\{a_{c}^{1},a_{c}^{2},\ldots,a_{c}^{n}\}$ where $a_{c}^{i}$ represents the $i^{th}$ audio frame in one data sample.
 Each audio frame will be converted into a vector representation: $e_{ac}^{i}=\text{Wav2Vec2}(a_{c}^{i})$.
@@
 In detail, the MHSA calculation process is as follows:
-$\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$ (1)
-$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (2)
+(3) $\text{Multihead}=\text{Concat}\text{(head}_{1},....,\text{head}_{h})W^{o}$
+(4) $\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$
 where $Q$ (queries) and $K$ (keys) of dimension $d_{k}$ and $V$ values of dimension $d_{v}$.
@@
 The attention scores are normalized using the softmax function:
-$\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$ (3)
+(5) $\text{Attention}(Q,K,V)=softmax(\frac{KQ^{T}}{\sqrt{d_{k}}})V$
 The attention function on a set of queries is calculated simultaneously packed together in a matrix Q.
@@
 Combining (2)-(4), this results in a matrix:
-$T_{ac}=\text{MHSA}(E_{ac})$ (4)
+(6) $T_{ac}=\text{MHSA}(E_{ac})$
 where $T_{ac}=\{t_{ac}^{1},t_{ac}^{2},\ldots,t_{ac}^{n}\}$ with size 520 $\times$ 512.
 $T_{ac}$ is then subjected to an average pooling layer to produce $T_{a}$, a condensed audio feature vector of size 512.
-3.2 Transcript Encoding
-The transcription process mirrors Audio Encoding, using SimCSE Gao et al.
-(2021)to extract sentence-level vector representations from earnings conference transcripts.
+4.2. Transcript Encoding
+The transcript encoding process mirrors Audio Encoding, using SimCSE (Gao et al., 2021) to extract sentence-level vector representations from earnings conference transcripts.
 SimCSE is a Siamese neural network architecture that learns to embed pairs of sentences into a shared space where similar sentences are mapped close together and dissimilar sentences are mapped far apart.
@@
 Then,$T_{tc}$ is subjected to the average pooling layer to produce $T_{t}$, where $T_{t}$ denotes the resultant extracted textual feature of size 768.
-This streamlined process effectively captures the textual nuances required for in-depth analysis.
-3.3 Earnings Conference Call Focuses Extraction And Analysis
-To obtain deep insights from an Earnings Conference Call on how it might influence future market performance, our approach encompasses several steps: (1) summarize ECCs, (2) extract investors’ focus information (3) calculate focus sentiment and (4) extract audio’s features of corresponding focus.
-Refer to the Figure 2 for additional details.
-(1) Summarize ECCs.
-To accommodate the token limitations of LLMs, we start by dividing the entire document into several chunks and summarizing each one.
-We then summary these summaries to create a comprehensive overview of the entire document.
-This two-tiered approach ensures that the summary capture both the detailed and overall information.
-In further, we will use LLM to get the embedding with a size 1024: $T_{s}=\text{LLM}(summary+chunk\ summaries)$
-(2) Extract the investor’s focus using RAG.
-In this step, we attempt to use LLM to examine and extract focuses of interest to investors.
-Therefore, we start by identifying the core topics that analysts typically focus on.
-We consult with four finance experts, each possessing extensive experience in financial analysis.
-These experts are tasked with meticulously reviewing earnings conference calls across various sectors to pinpoint and summarize the key topics that are frequently discussed and also hold the greatest interest for investors.
-Their analysis aims to identify overlaps where common discussion topics align closely with investor concerns, highlighting areas of particular significance and interest in the investment community.
-Once we have established this focused database, our goal is to locate where these topics are mentioned during the earnings calls and systematically extract these segments.
-However, accurately extracting this information from the dense narrative of earnings conference calls can pose a challenge.
-Because LLM searches are based on similarity, comparing a single-word topic to an entire document presents a scale issue, and a single word may lack the necessary context, thus diminishing the extraction capability.
-To enhance the precision of information extraction, we formulate multiple questions related to each topic but phrased differently, ensuring a broad coverage and increasing the likelihood of accurately pinpointing relevant sections.
-For instance, regarding the dividends topic, we pose questions such as: (1) "Have there been any changes in the stock dividends, and if so, at what rate have the dividends increased or decreased?" (2) Does the company Board expect to increase the stock dividends in the next future?
-(3) How does the dividend yield compare to the peers?
-After locating the segments where these topics are discussed, we extract the corresponding paragraphs.
-These paragraphs, however, may still contain irrelevant text such as preceding discussions.
-To address this, we use a contextual compression method from LLM, adept at distilling the paragraph down to its most relevant sentences, effectively obtaining the crucial data from the irrelevant text.
-(3) Calculate focus sentiment.
-Sentiment is a strong indicator for reflecting market perception.
-We determined to calculate the sentiment score based on the focus.
-Because this offers a more targeted and insightful view, which may be overlooked in a broader context.
-It provides a clearer picture of how specific developments or concerns are influencing investor behavior and market movements.
-(4) Extract audio’s features using RAG.
-Once we identify the focal points of interest, we aim to enrich our analysis by locating and examining the audio segments that correspond to these specific topics.
-By analyzing the audio features related to a particular focus, such as the company’s current projects, we can delve deeper into the emotional nuances and confidence levels exhibited by the speakers.
-This approach allows us to capture subtle cues in tone, pace, and emphasis that might indicate underlying sentiments or confidence about the discussed topics.
-The goal is to provide a more layered understanding of how speakers convey their messages and the potential impacts these emotional expressions have on the listeners’ perceptions and reactions.
-We utilize Praat Boersma and Van Heuven (2001) to extract vocal features, such as pitch, intensity, jitter, HNR(Harmonic to Noise Ratio) and etc, from ECC audio files.
-After obtaining the four key components, we then explore effective methods to merge these elements into a cohesive analysis.
-In further, we will use LLM to get the embedding with size 1024: $T_{f}=\text{LLM}(focus\ analysis+focus\ sentiment+focus\ audio\ features)$
-3.4 Additive Multi-modal Fusion
+4.3. Fine-grained Information Extraction from the Transcript by Using LLM
+To obtain deep insights from an Earnings Conference Call on how it might predict future market performance, our approach encompasses two parts: (1) summarize the transcript, and (2) important sentences extracted by using RAG.
+We show this process in Figure 2.
+We list all Prompts in Appendix A.1.
+4.3.1. Summarize the transcript
+In order to effectively summarize key information in lengthy ECC records, we employ a hierarchical summarization strategy.
+First, the entire document is divided into chunks, and then we use LLM to summarize each chunk individually.
+These individual summaries are then further summarized through LLM, resulting in a comprehensive summarization paragraph of the entire document.
+This two-layer approach ensures that both detailed and aggregate information is captured.
+In further, we use the OpenAI ‘text-embedding-3-small’ text embedding model to generate embeddings $T_{s}$ with size 1024 for summarized paragraph:
+(7) $T_{s}=\text{Embedding}(summary+chunk\ summaries)$
+4.3.2. Important sentences extracted by using RAG
+In this step, we aim to extract the most critical sentences containing relevant information from the entire transcript.
+To achieve this, we first convert the processed chunks into vector representations using the embedding model.
+Then, with input from financial experts, we design a list of questions about the ECC, which are stored in a ’Question Bank’ as a query list (see Appendix  A.2).
+For each question in the ’Question Bank’, we retrieve the relevant chunks from the vectors individually.
+Next, we use the context compressor, in combination with the LLM chat, to quiz each extracted chunk and query on its relevance.
+This process filters out only the sentences useful for answering the questions.
+Finally, we synthesize the query-selected sentences into a coherent prompt, to which the large language model will generate responses as the final answers to the queries.
+These generated answers are considered as the extracted important sentences.
+We then concatenate all the generated sentences and feed them into a text embedding model to generate the text embedding $T_{f}$ with size 1024:
+(8) $T_{f}=\text{Embedding}(\text{Concatenated}[selected\ sentences;])$
+4.4. Multimodal Fusion and Prediction
 Given the model’s reliance on several inputs and diverse data types, we identify an effective fusion structure to integrate these features into the training process to ensure a balanced weighting among components.
 We use additive interactions to handle the representational fusion of different abstract representations.
-These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline Liang et al.
-(2022).
+These operators can be viewed as differentiable building blocks that combine information from several different data streams and can be flexibly inserted into almost any unimodal pipeline (Liang et al., 2022).
 Given the audio feature $T_{a}$, textual feature $T_{t}$ from the transcript, and $T_{s},T_{f}$ from ECC analyzed text, additive fusion can be seen as learning a new joint representation:
-$E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{s}+w_{4}\cdot T_{f}+\epsilon$ (5)
+(9) $E=w_{0}+w_{1}\cdot T_{a}+w_{2}\cdot T_{t}+w_{3}\cdot T_{s}+w_{4}\cdot T_{f}+\epsilon$
 where $w_{1}\in R^{512\times 512}$, $w_{2}\in R^{768\times 512}$ and $w_{3},w_{4}\in R^{1024\times 512}$ are the weights learned for additive fusion, $w_{0}$ the bias term and $\epsilon$ the error term.
-$E$ is a vector with 512 as the final feature from the Earning Conference Call Encoder.
-3.5 Multi-Task Prediction
-We begin our prediction process by aggregating features from various modules into a comprehensive feature representation.
-This unified feature set is fed into a two-layer neural network designed to perform the regression task.
-Integrating these diverse inputs into a cohesive output is crucial, as it harnesses the strengths of each module to enhance analysis and prediction accuracy.
-Building on insights from previous research in multimodal financial risk prediction, which has demonstrated substantial improvements in prediction performance through multitask learning, we adopt a joint modeling approach.
-Here, we concurrently model volatility prediction and VaR prediction using a multi-task framework.
-The multi-task prediction module is comprised of two separate single-layer feedforward networks, each responsible for predicting volatility (vol) and Value at Risk (var) values individually.
-We train ECC Analyzer by optimizing multitask loss:
-$\mathcal{L}$ $=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})$ (6)
-$+(1-\mu)\text{max}(q\times(v-\hat{v}),(1-q)(\hat{v}-v))$
-multi-task learning allows us to optimize performance by accurately capturing and predicting multiple stock performance factors simultaneously.
-Table 1: Performance results on our proposed framework ECC Analyzer from different baseline models.
-Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$ $VaR$ Multi-Task
-Classical Method 0.713 1.710 0.526 0.330 0.284 / $\otimes$
-LSTM 0.746 1.970 0.459 0.320 0.235 / $\otimes$
-MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233 / $\otimes$
-HAN 0.598 1.426 0.461 0.308 0.198 / $\otimes$
-MRDM 0.577 1.371 0.420 0.300 0.217 / $\otimes$
-HTML 0.401 0.845 0.349 0.251 0.158 / D
-GPT-4-Turbo 2.198 3.187 5.059 7.959 11.824 0.371 D
-ECC Analyzer 0.316 0.553 0.306 0.247 0.159 0.049 D
-4 Results and Discussions
-4.1 Dataset
-The dataset utilized in this study is sourced from the publicly available S&P 500 ECC dataset as constructed by Qin and Yang (2019).
+$E$ is a vector with 512 as the final feature.
+This unified feature set $E$ is fed into two fully connected layers to perform the regression task.
+We train ECC Analyzer by optimizing the MSE loss:
+(10) $\mathcal{L}$ $=\mu(\sum_{i}(\hat{y_{i}}-y_{i})^{2})$
+5. Experiment
+We describe the datasets used in the experiment and the baseline settings for performance comparisons.
+Our experiments aim to answer the following research questions (RQs): RQ1: Can our proposed pipeline, when integrated with LLMs, significantly improve volatility prediction performance compared to current state-of-the-art (SOTA) approaches?
+RQ2: Does the synergy between our model and the LLM result in better performance compared to raw LLM predictions?
+RQ3: Does each distinct component derived from ECC contribute to improving volatility prediction accuracy?
+5.1. Dataset
+The dataset utilized in this study is sourced from the publicly available S&P 500 ECC dataset as constructed by (Qin and Yang, 2019).
 It includes both audio recordings and corresponding text transcripts from the 2017 earnings calls of 500 major companies listed on the S&P 500 and traded on U.S.
@@
 The dataset consists of 572 unique instances where the audio recordings were accurately and closely aligned with the text transcripts.
-Following the setup by Qin and Yang (2019), we partitioned the dataset into a training set and a test set with an 8:2 ratio, organized temporally to ensure that the data in the training set precedes those in the test set.
+Following the setup by (Qin and Yang, 2019), we partitioned the dataset into a training set and a test set with an 8:2 ratio, organized temporally to ensure that the data in the training set precedes those in the test set.
 This temporal division is crucial for maintaining the integrity of our predictive model, aligning the training process with the principle of using historical data to predict future risks—thus enhancing the accuracy and reliability of our forecasting approach.
-4.2 Baseline
-We compare our approach to several important baselines including 1) GARCH-based Classical Methods; 2) LSTM Gers et al.
-(2000) model; 3) MT-LSTM+ATT Luong et al.
-(2015) employing an attention-enhanced LSTM as the foundational model; 4) HAN (Glove) uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels for prediction; 5) MRDM Qin and Yang (2019) proposed a multi-modal deep regression approach for volatility prediction tasks; 6) HTML Yang et al.
-(2020) presented a state-of-the-art method; and 7) GPT-4-turbo-2024-04-09: directly utilize LLM to predict stock performance.
-We explain each baseline method in detail in the Appendix A.1
-4.3 Implementation Details
-We use GPT-4 for the ECC Analyzer experiment, utilizing it to analyze ECC data and build the analysis database.
-Throughout this process, we set the temperature parameter to 0.
-This ensures that the Large Language Models (LLMs) produce the most predictable responses, which aids in maintaining consistency in our experiments.
+5.2. Baseline Setup
+We compare our approach to several important baselines including:
+  • Classical Method: We incorporate the GARCH model and its derivatives, as described in  (Franses and Van Dijk, 1996; Kim and Won, 2018).
+This model is well-recognized for short-term volatility prediction but may not be as effective for forecasting average volatility over longer periods, such as n-day volatility.
+  • LSTM (Gers et al., 2000): The Long Short-Term Memory (LSTM) based method is a popular choice for financial time series prediction due to its efficacy in handling sequential data.
+We use a straightforward LSTM model as a benchmark for volatility prediction.
+  • MT-LSTM-ATT (Luong et al., 2015): combines the prediction of average n-day volatility with the forecasting of single-day volatility, employing an attention-enhanced LSTM as the foundational model.
+  • HAN (Glove): uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels.
+HAN first gets word embeddings using pre-trained Glove vectors and then processed by a Bi-GRU (Chung et al., 2014) encoder, while another Bi-GRU encoder simultaneously forms a sentence-level representation of each document.
+The resulting document representation is input into a regression layer to produce predictions.
+  • MRDM (Qin and Yang, 2019): The MRDM model first introduced a multi-modal deep regression approach to fuse the GloVe embeddings and hand-crafted acoustic features for volatility prediction tasks.
+  • HTML (Yang et al., 2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding.
+Similar to MDRM, HTML also leverages the same audio features.
+These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
+  • AMA-LSTM (Wang et al., 2024): The paper uses a strategy of adversarial training to minimize the effect of speaker gender in earnings conference calls.
+  • GPT-4-turbo-2024-04-09: We assessed the capability of LLMs in directly predicting volatility performance from ECCs.
+The model was set to generate a response with a zero temperature setting to ensure deterministic output.
+5.3. Implementation Details
+In the experiment, all interactions with LLMs were conducted using “GPT-4 Turbo-2024-04-09”.
+We set the temperature parameter to 0 to ensure that the LLMs produce the most predictable responses, thereby maintaining consistency in our experiments.
 For the overall training of the ECC Analyer framework, we developed the code using PyTorch.
-Each Multi-Head Attention layer in the network comprises 6 layers and 8 individual heads in each layer.
-The training process utilized batch sizes $b\in\{2,4,8\}$.
+Each Multi-Head Self-Attention layer in the model comprises 6 layers and 8 individual heads in each layer.
+The training process utilized batch sizes $b\in\{2,4,8,16\}$.
 We use a grid search to determine the optimal parameters and select the learning rate $\lambda$ for Adam optimizer among $\{1e-3,1e-5,1e-6,1e-7\}$.
-The best hyper-parameters were kept consistent across all experiments, with the exception of the trade-off parameter $\mu$ which varied between the two tasks.
-We list the evaluation metrics in Appendix A.2.
-4.4 Overall Results Analysis (RQ1)
-Table 1 shows the performance of various methods in predicting stock performance.
-Notably, the ECC Analyzer framework excels, especially in short-term and medium-term forecasts, with the lowest Mean Squared Error (MSE) values.
+5.4. Overall Results Analysis (RQ1 & RQ2)
+Table 1. Performance results on our proposed framework ECC Analyzer with different baseline methods.
+Model $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$
+Classical Method 0.713 1.710 0.526 0.330 0.284
+LSTM 0.746 1.970 0.459 0.320 0.235
+MT-LSTM-ATT 0.739 1.983 0.435 0.304 0.233
+HAN (Glove) 0.598 1.426 0.461 0.308 0.198
+MRDM 0.577 1.371 0.420 0.300 0.217
+HTML 0.401 0.845 0.349 0.251 0.158
+AMA-LSTM / 0.680 0.360 0.230 /
+GPT-4-Turbo 2.198 3.187 5.059 7.959 11.824
+ECC Analyzer 0.314 0.553 0.306 0.237 0.158
+Table 2.
+Performance results of ablation study.
+We designed the ablation study as follows: 1) Audio+Text: uses raw audio and text data from ECCs; 2)Audio+Text+$E_{os}$: adds an overall ECC summary generated by LLMs; 3)Audio+Text+$E_{cs}$: adds the chucks summary generated by LLMs; 4) Audio+Text+$E_{os}+E_{cs}$: integrates both overall and chunk summaries for the ECC; 5) Audio+Text+$E_{fo}$: combines raw data with focused analytical results; 6) $E_{os}+E_{cs}+E_{fo}$: merges all LLM analyses for prediction without raw data; 6) Audio+Text+$E_{os}+E_{cs}+E_{fo}$: combines all data and analyses for enhanced prediction.
+Module $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$
+Audio+Text 0.373 0.645 0.362 0.280 0.204
+Audio+Text+$E_{os}$ 0.373 0.638 0.380 0.276 0.201
+Audio+Text+$E_{cs}$ 0.375 0.640 0.385 0.275 0.201
+Audio+Text+$E_{os}+E_{cs}$ 0.357 0.627 0.335 0.267 0.199
+Audio+Text+$E_{fo}$ 0.324 0.579 0.323 0.230 0.165
+$E_{os}+E_{cs}+E_{fo}$ 0.343 0.601 0.344 0.247 0.179
+Audio+Text+$E_{os}+E_{cs}+E_{fo}$ 0.314 0.553 0.306 0.237 0.158
+Table 1 shows the performance of various methods in predicting volatility performance.
+Notably, the ECC Analyzer framework excels, especially in short-term forecasts (day 3 and day 7), with the lowest Mean Squared Error (MSE) values of 0.553 and 0.306, respectively.
 Its long-term prediction performance is comparable to the state-of-the-art method, HTML.
-This improvement highlights the benefits of deep analysis using LLMs extracted from ECCs.
-However, directly applying LLMs to stock performance prediction proves largely ineffective, similar to random guesses.
+However, the ECC Analyzer performs slightly lower than AMA-LSTM in predicting medium-term (day 15) volatility.
+Nevertheless, ECC Analyzer’s overall average MSE achieves superior performance.
+These encouraging experimental results illustrate that extracting fine-grained information from ECC data using LLMs can significantly enhance the model’s volatility prediction accuracy.
+In particular, the extracted fine-grained information features are especially beneficial for improving short-term forecasting performance, which is crucial for investors.
+In summary, addressing RQ1, the prediction performance of our proposed method surpasses that of the current SOTA approaches.
+In response to RQ2, directly applying LLMs to volatility prediction proves largely ineffective, akin to random guessing.
+These results also raise concerns about the potential misuse of LLMs, considering their impact on public safety.
+If investors use LLMs inappropriately for numerical outputs, they may increase financial risk.
 This indicates that LLMs are more effective as tools to enhance investors’ understanding of a company’s financial health rather than direct predictors of financial metrics.
-Additionally, the ECC Analyzer demonstrates outstanding performance in Value at Risk (VaR) prediction, confirming the efficacy of our methodology in providing a nuanced and comprehensive approach to financial risk prediction.
-We also report the results on Returns in Appendix B.
-Table 2: Performance results of ablation study.
-We designed the ablation study as follows: 1) Audio+Text: uses raw audio and text data from ECCs; 2)Audio+Text+$E_{os}$: adds a comprehensive ECC summary generated by LLMs; 3) Audio+Text+$E_{os}+E_{cs}$: integrates both overall and chunk summaries for the ECC; 4) Audio+Text+$E_{fo}$: combines raw data with focused analytical results; 5) $E_{os}+E_{cs}+E_{fo}$: merges all LLM analyses for prediction without raw data; 6) Audio+Text+$E_{os}+E_{cs}+E_{fo}$: combines all data and analyses for enhanced prediction.
-Module $\overline{MSE}$ $MSE_{3}$ $MSE_{7}$ $MSE_{15}$ $MSE_{30}$
-Audio+Text 0.373 0.645 0.362 0.28 0.204
-Audio+Text+$E_{os}$ 0.373 0.638 0.380 0.276 0.201
-Audio+Text+$E_{os}+E_{cs}$ 0.357 0.627 0.335 0.267 0.199
-Audio+Text+$E_{fo}$ 0.324 0.579 0.323 0.23 0.165
-$E_{os}+E_{cs}+E_{fo}$ 0.343 0.601 0.344 0.247 0.179
-Audio+Text+$E_{os}+E_{cs}+E_{fo}$ 0.310 0.553 0.306 0.22 0.159
-4.5 Ablation Study (RQ2)
+5.5. Ablation Study (RQ3)
 In our research, we conducted an ablation study to assess how different combinations of ECC analysis results impact our model’s performance.
 This systematic comparison helped us identify the individual contributions of each component.
+Each component is represented as follows: $E_{os}$ represents the embedding of the overall ECC summary; $E_{cs}$ represents the embedding of the summary for chunks; $E_{cs}$ represents the embedding of the extracted fine-grained important sentence.
 According to Table 2, we can find that the audio and text features extracted by advanced large language models significantly improved short-term prediction accuracy compared to previous methods.
@@
 Our best analytical results come from integrating the full spectrum of data and analytical outputs, underscoring the value of each component in our model.
-Notably, good predictive results were also can obtained using only analytics derived from LLMs, affirming the response to our RQ2: comprehensive analysis indeed enhances the predictive capability for stock performance.
+We also obtain good predictive results using only analytics derived from LLMs, affirming the response to our RQ3: comprehensive analysis indeed enhances the predictive capability for stock volatility prediction performance.
 Our findings also suggest that while earnings calls are information-rich, including every detail in the analysis can be counterproductive and may cloud essential insights.
 It is therefore critical to pinpoint and concentrate on the most predictive elements of the data, filtering out less relevant information to optimize the analysis process for stock performance prediction.
-4.6 Interpretability Study (RQ3)
-In our study, we initiate the analysis by collaborating with consultants to summarize and identify key focus points from earnings conference calls; the more comprehensive the summary, the better.
-Following this initial step, we combine these focus points and try different combinations to determine the optimal subset that most strongly predicts stock performance.
-(The finalized list of focus points is included in the appendix.) This method serves a dual purpose for investors.
-Firstly, it reveals which focus points are highly relevant to stock performance, allowing investors to prioritize their attention effectively.
-Secondly, by extracting and analyzing these key focuses, investors can understand the specific contents in the ECC that drive movements in the stock market.
-5 Conclusions
-This study confirms the effectiveness of using Large Language Models (LLMs) to extract and analyze key topics from earnings conference calls.
-Our approach not only pinpoints critical discussions but also assesses their impact on stock performance, predicting metrics like volatility and Value at Risk (VaR).
-By integrating textual and audio data, our model offers a comprehensive view, capturing subtleties such as tone and pitch.
-Results from benchmark comparisons demonstrate our model’s superior accuracy and predictive capabilities, highlighting LLMs’ potential to improve interpretability and decision-making in investments.
-Appendix A Supplement Material for Experiment Setup
-A.1 Baseline Setup
-  • Classical Methods: We incorporate the GARCH model and its derivatives, as described in  Franses and Van Dijk (1996); Kim and Won (2018).
-These models are well-recognized for short-term volatility prediction but may not be as effective for forecasting average volatility over longer periods, such as n-day volatility.
-  • LSTM Gers et al.
-(2000): Long Short-Term Memory Networks (LSTMs) are a popular choice for financial time series prediction due to their efficacy in handling sequential data.
-We use a straightforward LSTM model as a benchmark for volatility prediction.
-  • MT-LSTM+ATT Luong et al. (2015): combines the prediction of average n-day volatility with the forecasting of single-day volatility, employing an attention-enhanced LSTM as the foundational model.
-  • HAN (Glove): uses a Hierarchical Attention Network with dual-layered attention at the word and sentence levels.
-HAN first get word embeddings using pre-trained Glove vectors and then processed by a Bi-GRU Chung et al.
-(2014) encoder, while another Bi-GRU encoder simultaneously forms a sentence-level representation of each document.
-The resulting document representation is input into a regression layer to produce predictions.
-  • MRDM Qin and Yang (2019): The MRDM model first introduced a multi-modal deep regression approach to fuse the GloVe embeddings and hand-crafted acoustic features for volatility prediction tasks.
-  • HTML Yang et al.
-(2020): This work presented a state-of-the-art model that employs WWM-BERT for text token encoding.
-Similar to MDRM, HTML also leverages the same audio features.
-These unimodal features are then combined and processed through a sentence-level transformer, resulting in multimodal representations for each call.
-  • GPT-4-turbo-2024-04-09: We assessed the capability of LLMs in directly predicting stock performance from ECCs.
-The model was set to generate response with a zero temperature setting to ensure deterministic output.
-A.2 Evalation Metrics
-A.3 Volatility of Different Future Intervals
-We used the following formula to calculate the Volatility of Different Terms:
-$\sigma_{ijk}^{2}=\frac{1}{k-1}\sum_{m=1}^{k}[S_{i,(j-m)}-\frac{1}{k}\sum_{l=1}^{k}S_{i,(j-l)}]^{2}$
-Here, we define the $\sigma_{ijk}$ as the $k-days$ volatility of Stock $i$ at time $j$, which is calculated as the sample standard deviation of the past $k-days$ stock closed price of company $i$.
-But in the real application, we took the log form of historical volatility.
-A.4 VaR of Different Future Intervals
-Our second task is predicting the 1-day VaR of the target stock based on the multi-source inputs. The definition of VaR is:
-$x=F^{-1}(p)$
-The $F(\cdot)$ is the cumulative loss distribution, $p$ is the percentile we set, and $x$ is the VaR. From the idea of Quantile Regression, we can have:
-$L_{\tau}(y,\hat{y})=\begin{cases}\tau\cdot(y-\hat{y})&\text{if }y\geq\hat{y}\\ (1-\tau)\cdot(\hat{y}-y)&\text{if }y<\hat{y}\end{cases}$
-Calculating and estimating VaR can help the company better deal with financial risks and avoid extreme scenarios in the future.
-A.5 Returns with Different Future Intervals
-We used the following formula to calculate the Returns with Different Gaps:
-$R_{ijk}=\frac{S_{i,(j+k)}-S_{i,j}}{S_{i,j}}$
-Here, we denote $S_{i,j}$ as the closed price of company $i$’s stock at time $j$, and $R_{ijK}$ as the $k-day$ return of stock price $i$ at time $j$.
-We calculated this value as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$.
-Appendix B The Results on Returns
-To evaluate the predictive capability of our model, we have applied it to forecast stock returns. The results of these forecasts are systematically compiled and displayed in a table 3.
-Table 3: ECC Analyzer’s Prediction Performance On Stock Returns
-Return_3d Return_7d Return_15d Return_30d
-Predited Return 0.0024 0.0068 0.0131 0.0185
-True Return 0.0007 0.0017 0.0035 0.0139
-Error 0.0018 0.0051 0.0096 0.0046
-Percentage Error 2.6923 3.0964 2.7216 0.3348
-Appendix C Regression and Sentiment Analysis
-In the main chapter, we used the various returns and volatility of different terms as our predicted target.
-This part will introduce how we calculated these values and further results about the relationships between sentiments calculated by LLM and these target values.
-C.1 Cumulative Abnormal Returns and Regression Analysis
-We used both Abnormal Returns collected from the market and the returns of different terms to capture the sentiment delivered by ECCs.
-For the abnormal Returns, we used the following regression model Lee and Connolly (2010):
-$R_{it}=\alpha+\beta R_{mt}+\varepsilon_{it}$
-$R_{it}$ is the return of the security at time $t$, also $R_{mt}$ is the market portfolio return at time $t$.
-The Abnormal Returns we defined here are the difference between the real value of Returns and estimated ones, which is $\varepsilon_{it}$.
-For other dependent variables, we used the $R_{ijk}$ for $k=1,3,7,15,30$.
-Results are shown in Table 4.
-Table 4: This table shows the linear regression results of different kinds of returns.
-We used different types of returns as dependent variables.
-$CARs$ is the cumulative abnormal returns, calculated as the difference between the actual value of the predicted valve from factor models.
-The rest of the dependent variables are calculated as the difference between the closed price of company $i$ at time $j$ and time $j+k$ divided by the closed price of company $i$ at time $j$.
-The independent variable is the Sentiment score, which is calculated by LLM.
-We extracted all the events in one ECC and evaluated their sentiment score.
-The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
-Dependent Variable
-CARs Return_1d Return_3d Return_7d Return_15d Return_30d
-Constant $-0.0105^{***}$ $-0.0105^{***}$ $-0.0126^{***}$ $-0.0182^{***}$ $-0.0286^{***}$ $-0.0244^{***}$
-(0.004) (0.004) (0.004) (0.005) (0.007 (0.008)
-Sentiment $0.0223^{***}$ $0.024^{***}$ $0.0295^{***}$ $0.0442^{***}$ $0.0716^{***}$ $0.0852^{***}$
-(0.008) (0.008) (0.009) (0.010) (0.015) (0.016)
-Adjusted $R^{2}$ 0.016 0.018 0.022 0.038 0.048 0.055
-Observation 572 572 572 572 572 572
-For the volatility, we used the difference between the two closed periods of volatility.
-The dependent variable for volatility could be calculated as $log(\sigma_{i(j+1)k})-log(\sigma_{ijk})$ for $k=3,7,15,30$.
-Results are shown in Table 5.
-Table 5: This table shows the linear regression results of different terms of Volatility.
-We used four different terms of volatility.
-Volatility here is defined by the log form of the sample standard deviation of the past $k-days$ stock closed price of company $i$.
-The independent variable is the Sentiment score, which is calculated by LLM.
-We extracted all the events in one ECC and evaluated their sentiment score.
-The percentage of events with positive sentiment minus the percentage of events with negative sentiments calculates the score.
-Dependent Variable
-Volatility_3d Volatility_7d Volatility_15d Volatility_30d
-Constant $-0.0254^{***}$ $-0.017^{***}$ $-0.0071^{***}$ $-0.005^{***}$
-(0.013) (0.009) (0.006) (0.004)
-Sentiment $-0.0681^{***}$ $-0.0556^{***}$ $-0.0438^{***}$ $-0.0266^{***}$
-(0.026) (0.018) (0.013) (0.009)
-Adjusted $R^{2}$ 0.013 0.019 0.022 0.019
-Observation 572 572 572 572
-From Table 4 and Table 5, we can find that the Sentiment generated from the LLM is always statistically significant in the linear regression model results.
-Comparing the coefficient of different terms of Returns, we found that the coefficient gets larger when there are more significant term gaps.
-This indicates that the impact of the ECC’s sentiment will appear as time progresses.
-There is a higher probability that the impact of the ECCs occurs after a certain period because, besides the financial data, most information, like future projects mentioned in ECCs, will not be reflated instantly.
-That could explain why the coefficient will rise as the period gets larger.
-However, the volatility change will have a smaller impact when the period gets larger(because the absolute value of Sentiments’ coefficient in Table 5 is smaller when the period gets more extensive), which means that the same ECCs sentiments effect will have a more significant impact in short terms rather than the longer terms.
-Figure 3: We asked four experts to help us analyze the possible scenarios of ECCs.
-Our experts focus on the following general Categories of the ECCs’ transcript: Financial Indicators, Employee & Managers, Cost, Project Expansion, Business, and Future Look.
-Each general category is separated into some minor focusing points.
-Our experts put forward three questions for each small focusing point.
-We used these questions to ask LLM and wished for a directional response in the thinking pattern hidden in the questions that we wanted to teach the LLM.
-Appendix D Prompt Design
-D.1 Prompt for Summarizing Earnings Conference Call Segments
+6. Conclusion
+In this paper, we propose the ECC Analyzer, a novel framework that leverages large language models (LLMs) for in-depth analysis of ECC data to enhance volatility prediction performance.
+The ECC Analyzer extracts information from ECCs at multiple levels to assist in volatility prediction.
+Our experiments demonstrate that our proposed method improves the overall average Mean Squared Error (MSE).
+Specifically, we achieve better results in short-term forecasting, with medium-term and long-term forecasts also on par with SOTA methods.
+Additionally, we analyze the role of each component in the proposed framework for the forecasting process through ablation experiments, highlighting the contribution of each element to the overall performance.
+Appendix A Appendix
+A.1. Prompt Design
+A.1.1. Prompt for Summarizing Earnings Conference Call Segments
   • Identify Key Points
@@
     If applicable, relate the segment’s content to broader company objectives or industry trends to provide context and show how the segment fits into the bigger picture.
-D.2 Prompt for Creating an Overview Summary from Earnings Conference Call Segments
+A.1.2. Prompt for Creating an Overview Summary from Earnings Conference Call Segments
   • Gather Segment Summaries
@@
     Look for common themes, recurring issues, or consistent messages across the segments. Note any overarching strategies, goals, or concerns expressed by the company executives.
-D.3 Prompt for Extracting Focus And Explore Its Impact
-  • Listen to the Call
+A.1.3. Prompt for Extracting Important Sentences.
+  • Check to the Call
     Begin by thoroughly listening to the entire earnings conference call. Pay attention to both the prepared remarks and the question-and-answer session.
@@
     Identify statements or discussions that involve significant financial metrics, strategic initiatives, new products or markets, regulatory impacts, or any notable shifts in operations.
-These are potential focus points that could influence investor perceptions and stock price.
+These are potential focus points that could influence investor perceptions and stock prices.
   • Document Evidence
@@
   • Analyze Impact on Stock Movement
-    Pre and Post Analysis: Examine stock price movements immediately before and after the call to capture initial reactions.
+    Pre and Post-Analysis: Examine stock price movements immediately before and after the call to capture initial reactions.
   • Longer-term Impact
@@
     Ensure to factor in overall market conditions and sector movements to isolate the impact of the earnings call from broader market trends.
-D.4 Prompt for Analyzing Sentiment of Focus Points and Supporting Evidence from an Earnings Conference Call
-  • Analyze Sentiment for Each Focus Point
-    Apply the sentiment analysis to the text surrounding each focus point.
-Pay attention to the language used, such as positive, negative, or neutral descriptors, and the intensity of the language.
-Please assign a continuous decimal sentiment score to each event, ranging from -1 to 1.
-A score of -1 represents a highly negative sentiment, 0 indicates neutrality, and 1 signifies a highly positive sentiment.
-  • Context Consideration
-    Consider the context in which each point was discussed. Assess whether the sentiment is directly related to the focus point or influenced by broader discussion themes.
-  • Document Supporting Quotes
-    For each focus point, document specific quotes or statements from the call that illustrate the sentiment. Note the speaker and their role to add credibility to the sentiment analysis.
-D.5 Prompt for Analyzing Audio Features of Focus Points from an Earnings Conference Call
-  • Tone and Pitch Analysis
-    Examine variations in tone and pitch within each audio segment. Look for patterns that might indicate emphasis, uncertainty, confidence, or stress.
-  • Volume and Speech Rate
-    Measure changes in volume and variations in speech rate. High volume and rapid speech may indicate areas of strong emotion or importance.
-  • Pause Patterns
-    Identify the frequency and duration of pauses, which can provide insights into the speaker’s thought process or hesitation.
+A.2. Design of Question Bank
+Table 3. The designed Question Bank
+Focus Category Focus Item Questions
+Financial Indicator Dividend Q1: Did this company pay the investors dividend?
+Q2: Have there been any increases or decreases in the stock dividends?
+If yes, what is the rate at which the dividends have been increasing?
+Q3: What type of dividend did the company pay?
+Revenue Q1: What was the company’s reported revenue for the past quarter, and how does it compare to the same quarter in the previous year?
+Q2: What factors influenced the company’s revenue performance this quarter?
+Q3: What are the company’s revenue forecasts for the upcoming quarters, and what strategies are in place to achieve these targets?
+Return Q1: What was the company’s net profit margin for this quarter, and how has it changed from the previous quarter or year?
+Q2: What is the company’s Return on Equity (ROE) and Return on Assets (ROA) for this period?
+Q3: How does the company plan to enhance shareholder value in the upcoming periods?
+Are there any dividends or buybacks planned?
+Earnings Q1: Have there been any increases or decreases in the earnings?
+If yes, what is the rate at which the earnings have been increasing or decreasing?
+Q2: What is the outlook provided by the executives of this company in relation to the future earnings growth?
+Q3: Are the earnings above or below compared to the expectations?
+Are they attractive compared to the peers?
+Employee Manager Salary Q1: What percentage of the company’s total expenses is currently allocated to employee salaries, and how has this changed in response to recent business developments?
+Q2: How does your company’s compensation structure compare with industry standards, particularly in terms of salary, benefits, and bonuses?
+Q3: What were the average salary increases or decreases across the company this year compared to last year?
+Pension Q1: What is the current status of the company’s pension fund, and what were the major changes to its funding status over the past year?
+Q2: How does the company manage its pension liabilities, and what strategies are in place to address any underfunded positions?
+Q3: What are the expected impacts of current pension commitments on the company’s future financial performance?
+Management Change Q1: Have there been any recent changes in the company’s key management positions, including the CEO, etc.
+Q2: What were the reasons behind any recent management changes, particularly in the CEO position?
+Q3: What impact are the recent management changes expected to have on the company’s strategy in the near term?
+Cost Operating Costs Q1: What were the total operating costs this quarter compared to the previous quarter?
+Q2: Which factors contributed to any significant changes in operating costs?
+Q3: How are you managing operating costs in light of current economic conditions?
+Cost of Goods Sold Q1: How did the Cost of Goods Sold change this quarter, and what were the driving factors behind these changes?
+Q2: What percentage of revenue does the Cost of Goods Sold represent, and how does this compare to industry norms?
+Q3: Are there any initiatives in place to reduce Cost of Goods Sold without compromising quality?
+Marketing and Sales Costs Q1: How much did the company spend on marketing and sales this quarter?
+Q2: What specific marketing or sales strategies contributed to these costs?
+Q3: Are there plans to adjust these strategies in the upcoming quarters based on performance?
+Expansion Geographic Expansion Q1: What specific regions or markets is the company expanding into, and what factors influenced these selections?
+Q2: What are the initial costs associated with the geographic expansion and strategies are in place to support it?
+Q3: What is the projected timeline for the new regional operations to reach profitability?
+Product Line Expansion Q1: What new products is the company introducing, and what consumer or market needs do they aim to address?
+Q2: How will the introduction of these products impact the company’s production costs and overall financial performance?
+Q3: Are there any expected synergies between the new products and existing products or services?
+Market Segmentation Expansion Q1: Which new customer segments is the company targeting, and what research supports this strategic direction?
+Q2: What marketing strategies will be employed to reach these new segments, and what are the anticipated costs?
+Q3: What are the growth expectations for these new market segments over the next few years?
+Business Business Q1: What are the key projects or initiatives currently being undertaken by the company, and provide a brief overview of what each project entails?
+Q2: How has the performance of these projects compared to the previous quarter, and how do they stand in relation to competitors in the same sector?
+What has been the market’s response to these initiatives?
+Q3: What are the generated revenues from these projects for the current reporting period, and what potential risks could impact their future performance?
+Future Outlook Future Outlook Q1: What are the company’s primary strategic goals for the upcoming year, and what key initiatives are planned to achieve these objectives?
+Q2: How do these future plans align with current industry trends and market demands?
+Q3: What are the expected financial impacts of these plans on the company’s performance in the short and long term?